Cochrane, Inc., is considering a new three-year fixed expansion project that requires an initial fixed asset investment of $1,860,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $1,950,000 in annual sales, which costs of $1,060,000. If the tax rate is 35 percent.

In the previous problem, suppose the required return on the project is 14 percent. What is the project's NPV?

NVP

795500[1-(1.14)^-3]/0.14

795500*2.32163203

1846858.28-1,860,000

~$-13141.72

Suppose the project requires an initial investment in net working capital of $150,000, and the fixed asset will have a market value of $175,000 at the end of the project. What is the project's Year 0 net cash flow? Year 1? Year 2? Year 3? What is the new NPV?
Answer #1
Question a:
Particulars 0 1 2 3
Initial Investment
Purchase of Fixed Asset (A) -1860000
Operating Cash flows
Sales Revenue 1950000 1950000 1950000
Less: Annual Costs -1060000 -1060000 -1060000
Less: Depreciation
(1860000/3 years)
-620000 -620000 -620000
Profit before Tax 270000 270000 270000
Less: Tax @35% -94500 -94500 -94500
Profit After Tax 175500 175500 175500
Add back Depreciation 620000 620000 620000
Net Operating Cash flows (B) 795500 795500 795500
Total Cash flows (C = A+B) -1860000 795500 795500 795500
Discount Factor @14% (D)
1/(1+14%)^n
n=0,1,2,3
1 0.877193 0.769468 0.674972
Discounted Cash flows (E = C*D) -1860000 697807 612111.4 536939.8
NPV of the project -13141.72
Therefore, NPV of the project is -13141.72
Question b:
Particulars 0 1 2 3
Initial Investment
Purchase of Fixed Asset -1860000
Investment in Working Capital -150000
Net Investment (A) -2010000
Operating Cash flows
Sales Revenue 1950000 1950000 1950000
Less: Annual Costs -1060000 -1060000 -1060000
Less: Depreciation
(1860000/3 years)
-620000 -620000 -620000
Profit before Tax 270000 270000 270000
Less: Tax @35% -94500 -94500 -94500
Profit After Tax 175500 175500 175500
Add back Depreciation 620000 620000 620000
Net Operating Cash flows (B) 795500 795500 795500
Terminal Value
Salvage value pf asset 175000
Less: Tax@35% -61250
Net salvage value of machine 113750
Working capital recovered 150000
Net Terminal Value (C ) 263750
Total Cash flows (D = A+B+C) -2010000 795500 795500 1059250
Discount Factor @14% (E)
1/(1+14%)^n
n=0,1,2,3
1 0.877193 0.769468 0.674972
Discounted Cash flows (F = D*E) -2010000 697807 612111.4 714963.6
NPV of the project 14882.01
Year 0 Cash flow is -$2010000
Year 1 Cash flow is 697807
Year 2 Cash flow is 612111.4
Year 3 Cash flow is 714963.6
New NPV of the project is 14882.01